OpenStudy (anonymous):

Can someone be kind and help me please and thank you.... What is the factored form for 9x^4-6x^3+18x^2-12x 4x^3+3x^2+8x+6 20p^3+40p^2+15p+30

OpenStudy (jjuden):

for eachone?

OpenStudy (anonymous):

yes please

OpenStudy (jjuden):

ok

OpenStudy (jjuden):

first one: 3x(x^2+2)(3x-2)

OpenStudy (jjuden):

second one: (x^2+2)(4x+3)

OpenStudy (dtan5457):

@jjuden Can you explain them too?

OpenStudy (jdoe0001):

wondering what part of that is meant to be "help"

OpenStudy (jjuden):

hold on

OpenStudy (jdoe0001):

\(\bf 9x^4-6x^3+18x^2-12x\implies (9x^4-6x^3)+(18x^2-12x)\) take common factors on each, see what you get

OpenStudy (anonymous):

@jdoe0001 i know 3 is a common factor for each number but not sure how to do it

OpenStudy (anonymous):

@jjuden thanks I'm pretty sure you were going to explain afterwards

OpenStudy (jdoe0001):

well.. the 1st group has a common factor of \(\bf 3x^3\) the 2nd group has a common factof or \(\bf 6x\) try those

OpenStudy (anonymous):

is there some type of formula to solve these equations

OpenStudy (jdoe0001):

nope

OpenStudy (jdoe0001):

you'd have to look at what's common to the terms

OpenStudy (jdoe0001):

\(\bf 9x^4-6x^3+18x^2-12x\implies (9x^4-6x^3)+(18x^2-12x) \\ \quad \\ (3\cdot {\color{brown}{ 3\cdot xxx}}x-{\color{brown}{ 3}}\cdot 2\cdot {\color{brown}{ xxx}})+({\color{brown}{ 6}}\cdot 3\cdot {\color{brown}{ x}}x-{\color{brown}{ 6}}\cdot 2{\color{brown}{ x}})\) that should make them noticeable

OpenStudy (jdoe0001):

\(\bf 9x^4-6x^3+18x^2-12x\implies (9x^4-6x^3)+(18x^2-12x) \\ \quad \\ 3x^3{\color{blue}{ (3x-2)}}+6x{\color{blue}{ (3x-2)}}\impliedby \textit{notice any other common factor?}\)

OpenStudy (anonymous):

yes 3x and 2

OpenStudy (jdoe0001):

well.. notice is a binomial though thus \(\bf 9x^4-6x^3+18x^2-12x\implies (9x^4-6x^3)+(18x^2-12x) \\ \quad \\ 3x^3{\color{blue}{ (3x-2)}}+6x{\color{blue}{ (3x-2)}}\implies {\color{blue}{ (3x-2)}}(3x^3+6x) \\ \quad \\ (3x-2)({\color{brown}{ 3x}}xx+{\color{brown}{ 3}}\cdot 2{\color{brown}{ x}})\impliedby \textit{see any other common factor?}\)

OpenStudy (anonymous):

ok...i think i get it a little

OpenStudy (anonymous):

I'm sorry math is really not my strong suit

OpenStudy (jdoe0001):

well.. one of the binomials has a common factor that is, 3x so \(\bf 9x^4-6x^3+18x^2-12x\implies (9x^4-6x^3)+(18x^2-12x) \\ \quad \\ 3x^3{\color{blue}{ (3x-2)}}+6x{\color{blue}{ (3x-2)}}\implies {\color{blue}{ (3x-2)}}(3x^3+6x) \\ \quad \\ (3x-2)({\color{brown}{ 3x}}xx+{\color{brown}{ 3}}\cdot 2{\color{brown}{ x}})\implies (3x-2)3x(x^2+2)\)

OpenStudy (anonymous):

ohhhhhh ok ok ok...when you do it like that i understand better

OpenStudy (jdoe0001):

the 2nd one is about the same thing you group in 2 terms groups so \(\bf 4x^3+3x^2+8x+6\implies (4x^3+3x^2)+(8x+6)\) and get a common factor for each you'd end up with a common binomial \(\bf 4x^3+3x^2+8x+6\implies (4x^3+3x^2)+(8x+6) \\ \quad \\ x^2{\color{blue}{ (4x+3)}}+2{\color{blue}{ (4x+3)}}\implies {\color{blue}{ (4x+3)}}(x^2+2)\)

OpenStudy (jdoe0001):

and the last one is about the same, \(\bf 20p^3+40p^2+15p+30\implies (20p^3+40p^2)+(15p+30) \\ \quad \\ 20p^2{\color{blue}{ (p+2)}}+15{\color{blue}{ (p+2)}}\)

OpenStudy (anonymous):

so the common factor would be 20 and 15?

OpenStudy (jdoe0001):

for the groups respectively, yes

OpenStudy (jdoe0001):

if you multiply \(20p^2\) and distribute, you'd get the original 1st group

OpenStudy (anonymous):

jesus take the wheel

OpenStudy (anonymous):

5(4p^2+3)(p+2)???????

OpenStudy (anonymous):

no wait sorry...5(2p^2+6)(p+4)

OpenStudy (jdoe0001):

\(\bf 20p^3+40p^2+15p+30\implies (20p^3+40p^2)+(15p+30) \\ \quad \\ 20p^2{\color{blue}{ (p+2)}}+15{\color{blue}{ (p+2)}}\implies {\color{blue}{ (p+2)}}(20p^2+15) \\ \quad \\ (p+2)(4\cdot 5pp+3\cdot 5)\implies (p+2)5(4pp+3) \\ \quad \\ 5(4p^2+3)(p+2)\)

OpenStudy (anonymous):

dammit man

OpenStudy (anonymous):

thanks sooo much for your help @jdoe0001

OpenStudy (jdoe0001):

yw